# The Identification Power of

### v1

@@ 4 Bounds Analysis
An important parameter in this class is the distribution function of the treatment effect, $F_{Y_{1}-Y_{0}}^{\ast}(\delta)$ (Example 2.2), which corresponds to the choice $\varphi(y_{1},y_{0})=y_{1}-y_{0}$. As a special case, the fraction of positive treatment effect (Example 2.1) is given by $\mathbb{P}(Y_{1}>Y_{0})=1-F_{Y_{1}-Y_{0}}^{\ast}(0)$.
@@
where $\tilde{\varphi}_{y}(\delta|x)=\sup\left\{y_{0}\in\mathcal{Y}_{0}(x):\varphi(y,y_{0})<\delta\right\}$. Note that for a fixed $\delta$, the set $\left\{y_{0}\in\mathcal{Y}_{0}(x):\varphi(y,y_{0})<\delta\right\}$ for some $y\in\mathcal{Y}_{1}(x)$ and $x\in\mathcal{X}$ may be empty. In such cases, we define $\tilde{\varphi}_{y}(\delta|x)$ as minus infinity.
Fan et al. (2017) show that when $\psi$ is a $\varphi$-indicator function, the identified set $\widetilde{\Theta}_{I}$ based on $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$ corresponds to the interval $\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$, where $F_{\varphi}^{L}(\delta)=\mathbb{E}[F_{\min,\varphi}(\delta|X)]$ and $F_{\varphi}^{U}(\delta)=\mathbb{E}[F_{\max,\varphi}(\delta|X)]$. Hence, by Proposition 3.1, the identified set $\Theta_{I}$ from the experimental data is also given by $\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$. Fan et al. (2017) also examine the identification power of covariates $X$ through these analytical expressions for the sharp bounds.
We can similarly characterize the identified set $\widetilde{\Theta}{IC}$ based on the self-selection-joint CDFs $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$. Let $\mathcal{Y}_{1}(s,x)$ and $\mathcal{Y}_{0}(s,x)$ denote the supports of $Y_{1}$ and $Y_{0}$ given $(S,X)=(s,x)$, respectively. Then $\widetilde{\Theta}_{IC}$ corresponds to the interval $\left[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)\right]$, where $F_{\varphi}^{L}(\delta)=\mathbb{E}[F_{\min,\varphi}(\delta|S,X)]$ and $F_{\varphi}^{U}(\delta)=\mathbb{E}[F_{\max,\varphi}(\delta|S,X)]$ with
@@
Therefore, by Theorem 3.2, the identified set $\Theta_{IC}$ under the combined data corresponds to the interval $\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$ when $\theta_{o}$ is represented by a $\varphi$-indicator function. The key difference between the identified sets lies in the inclusion of the self-selection variable $S$ in the conditioning events in equations (11) and (12). The variable $S$ is expected to carry information about the dependence structure between $Y_{1}$ and $Y_{0}$, thereby tightening the identified set.
@@
Proposition 4.2.
@@
Proposition 4.2 (ii) is our novel contribution, providing analytical expressions for the sharp bounds of $\theta_{o}$—when represented by a $\varphi$-indicator function—under the combination of experimental and observational data. This result facilitates the computation of the identified set using the combined data.
@@
To illustrate the identifying power of the combined data, we present simple numerical examples in Appendix A. Focusing on $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$, these examples show that the identified set can shrink substantially when self-selection depends on the potential outcomes.
Similar to Theorem 4.2 for super-modular functions, we can establish sufficient and necessary conditions under which $\Theta_{IC}$ is a proper subset of $\Theta_{I}$. To simplify the technical argument, Theorem 4.5 below presents such a result for the case that $\mathcal{Y}_{d}(s,x)=\mathcal{Y}_{d}$ for $d=0,1$ and all $(s,x)\in\{0,1\}\times\mathcal{X}$.
Theorem 4.5.
Suppose that $\psi$ is a $\varphi$-indicator function with $\varphi$ being continuous and non-decreasing in each argument, and that the condition in Proposition 4.2 (ii) holds. Suppose further that $\mathcal{Y}_{d}(s,x)=\mathcal{Y}_{d}$ for $d=0,1$ and all $(s,x)\in\{0,1\}\times\mathcal{X}$. Then $\Theta_{IC}=\Theta_{I}$ if and only if there exist $\bar{y}$ and $\underline{y}$ such that $F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1]$ attains its maximum at $\bar{y}$ and minimum at $\underline{y}$ for almost all $(s,x)\in\{0,1\}\times\mathcal{X}$.
A sufficient condition for the “if and only if” statement in Theorem 4.5 is that the self-selection variable $S$ is independent of the potential outcomes $(Y_{1},Y_{0})$ conditional on $X$. This implies that, in general, the combined data can tighten the identified set when self-selection depends on the potential outcomes.
@@ 5 Additional Restrictions and Computational Approach
As in the philosophy of partial identification analysis (Manski, 2003), additional model restrictions can help further narrow the identified set for $\theta_{o}$ when they are plausible. However, such restrictions may also complicate the computation of the identified set, particularly by making it difficult to derive analytical expressions for the sharp bounds, as was done in the previous section. In this section, we introduce two additional restrictions that are plausible in many empirical contexts and present a computational approach to obtain the identified set under these restrictions.
@@ 5.1 Additional Restrictions
The first restriction we consider is mutual stochastic monotonicity of the potential outcomes, as studied by Frandsen and Lefgren (2021). We say that the potential outcomes $Y_{1}$ and $Y_{0}$ are mutually stochastically increasing if they satisfy the following condition.
Assumption 5.1 (Mutually Stochastically Increasing).
The conditional distributions $\mathbb{P}(Y_{1}\leq t\mid Y_{0}=y,S=s,X=x)$ and $\mathbb{P}(Y_{0}\leq t\mid Y_{1}=y,S=s,X=x)$ are each non-increasing in $y$ almost everywhere, for almost all $(s,x)\in{0,1}\times\mathcal{X}$. [^10] Conditioning on $S$ and $X$ may render this assumption restrictive. Alternatively, one may impose the same condition without conditioning on $S$ and $X$, resulting in a weaker assumption with less identifying power.
This assumption implies that individuals with higher potential outcomes under one treatment state are more likely to have higher potential outcomes under the other state. Such an assumption is plausible in many empirical contexts—for example, in a small-class-size program, students who perform well academically in either small or regular-sized classes are also likely to perform well in the alternative class size. Frandsen and Lefgren (2021) demonstrate that this assumption substantially tightens the identified set for the distribution of treatment effects. Chetty et al. (2017) also employ a similar restriction to sharpen bounds on DTE parameters in their empirical study of income mobility.
@@
This assumption implies that individuals who self-select into treatment are more likely to experience higher treatment effects than those who self-select into no-treatment. [^12] It may be more natural to formulate this assumption solely based on observational data by replacing the inequality $\mathbb{P}(Y_{1}-Y_{0}>c\mid S=1,X=x)\geq\mathbb{P}(Y_{1}-Y_{0}>c\mid S=0,X=x)$ in Assumption 5.2 with $\mathbb{P}(Y_{1}-Y_{0}>c\mid D=1,G=\mathrm{obs},X=x)\geq\mathbb{P}(Y_{1}-Y_{0}>c\mid D=0,G=\mathrm{obs},X=x)$. Under the maintained Assumptions 2.12.5, this version is equivalent to Assumption 5.2. We refer to this as the generalized Roy model selection assumption, as it is implied by the selection behavior in the generalized Roy model. This assumption pertains specifically to observational data, as it restricts the self-selection mechanism. Hence, a further benefit of incorporating observational data is the ability to impose such restrictions on the self-selection mechanism to improve identification.
@@ 5.2 Computational Approach
We propose a computational approach to obtain the sharp bounds for $\theta_{o}$ under the combined data and either or both of the additional Assumptions 5.1 and 5.2. When Assumptions 5.1 and 5.2 are imposed in addition to Assumptions 2.12.5, the sharp lower and upper bounds for $\theta_{o}$ under the combined data can be obtained by solving the following minimization and maximization problems:
@@
$F_{Y_{d}|Y_{d^{\prime}}SX}(t|y,s,x)\geq F_{Y_{d}|Y_{d^{\prime}}SX}(t|y^{\prime},s,x);$ (15)
$\mbox{ for almost all }(y,y^{\prime},d,d^{\prime},s,x)\mbox{\ with\ }y^{\prime}\geq y\mbox{ and }d\neq d^{\prime};$
@@
Constraint (14) is derived from Theorem 3.2, which shows that the identified set for $\theta_{o}$ is fully characterized by $F_{Y_{1}SX}^{\ast}$ and $F_{Y_{0}SX}^{\ast}$. Constraints (15) and (16) are imposed by Assumptions 5.1 and 5.2, respectively. Note that the optimization problem (13)–(16) is not formulated over the treatment variable $D$ or the data source indicator $G$, as constraint (14) already contains the identifying information conveyed by these variables. This formulation reduces computational cost by minimizing the number of optimization variables.
When all random variables are discrete, the optimization problem (13)–(16) can be formulated as a finite-dimensional linear program, to which efficient algorithms and solvers for linear programs can be applied. See Appendix B for details of the linear programming formulation. Inference methods for linear program solutions, including those proposed by Fang et al. (2023) and Cho and Russell (2024), can also be applied in this setting.
Let $\widetilde{\mathcal{F}}^{\dagger}$ denote the class of CDFs $F$ of all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5 and 5.15.2, with $F^{\ast}$ replaced by $F$. The identified set for $\theta_{o}$ under the combined data and these assumptions is defined as
$\Theta_{IC}^{\dagger}\equiv\big{\{}\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]:F\in\widetilde{\mathcal{F}}^{\ast}\big{\}}$
@@
The following proposition shows that $\Theta_{IC}^{\dagger}$ is an interval characterized by the solutions to the minimization and maximization problems (13)–(16).
@@
This result enables us to compute the identified set for $\theta_{o}$ under the additional restrictions (Assumptions 5.1 and 5.2) by solving the optimization problem (13)–(16). Although we focus on the case where both assumptions are imposed jointly, the sharp bounds can also be computed by solving a similar optimization problem when either assumption is imposed individually.
@@ 6 Empirical Illustration
We illustrate the identification result using the DRPT data from Gaines and Kuklinski (2011), conducted in Illinois as part of the Cooperative Campaign Analysis Project 2008. The study evaluated the impact of negative campaign advertisements on candidate evaluations during the 2008 U.S. presidential election.
The sample consists of 483 adult residents of Illinois, randomly assigned to three groups: treatment (118 residents), no-treatment (129 residents), and self-selection (236 residents). Individuals in the treatment group were exposed to negative campaign e-flyers about John McCain and Barack Obama, while those in the no-treatment group were not. Participants in the self-selection group chose whether to view the material, with 90 of 236 opting in. The outcome variable $Y$ is each respondent’s feeling thermometer rating toward each candidate, ranging from 0 (unfavorable) to 100 (favorable), with 50 representing a neutral stance. The dataset includes a single covariate $X$ indicating each respondent’s partisanship: Republican, Democrat, or Independent.
Let $Y_{d,\text{McCain}}$ and $Y_{d,\text{Obama}}$ denote the potential feeling thermometer ratings toward John McCain and Barack Obama, respectively, under treatment status $d\in\{0,1\}$, where $d=1$ indicates exposure to the negative campaign flyer. Our parameter of interest is $\mathbb{P}(Y_{1,j}<Y_{0,j})$, the proportion of individuals whose evaluation of candidate $j\in\{\text{McCain},\text{Obama}\}$ is negatively affected by the negative campaign material. A notable feature of this parameter is that it remains well-defined even if the outcome $Y_{d,j}$ is ordinal rather than cardinal. This is particularly relevant in our application, as the feeling thermometer rating reflects individuals’ subjective assessments and lacks strong cardinal meaning (see, e.g., Wilcox et al. (1989)). Many other common causal parameters such as the ATE, $\mathbb{E}[Y_{1,j}-Y_{0,j}]$, are less meaningful in this context.
@@
Table 1 presents the estimated identified sets of $\mathbb{P}(Y_{1}<Y_{0})$ for each candidate, for the full population and specified subpopulations, both with and without the sample from the self-selection group ($G=\mathrm{obs}$). These identified sets are estimated using Proposition 4.2 (with $\varphi(y_{1},y_{0})=y_{1}-y_{0}$ and $\delta=0$), in conjunction with Lemma 3.1, where the true distribution $F_{YDGX}^{\ast}$ of the observables is replaced by the empirical distribution of the observables.
The results show that the identified sets based solely on the random-assignment sample ($G=\mathrm{exp}$) are wide and not very informative. However, incorporating the self-selection sample ($G=\mathrm{obs}$) substantially tightens the estimated identified sets—by 22% for John McCain and 25% for Barack Obama. Focusing on Barack Obama, for instance, the results indicate that at least 23% of individuals are negatively affected by the negative campaign advertisement, while no less than 23% appear resistant to it. This suggests that the advertisement has a substantial impact, but it is not uniformly effective. In particular, for the subpopulation of the opt-in group ($S=1$), the estimates imply that at least 30% are negatively affected, while no less than 34% are unaffected.
These findings demonstrate the empirical relevance of combining random-assignment and self-selection data sources. In particular, they highlight a novel advantage of DRPT designs in improving the informativeness of distributional treatment effect or treatment effect heterogeneity analyses.

### v4

@@ 4.2 Identified Sets for $\varphi$-indicator Functions and the Identification Power of Data Combination
An important parameter in this class is the distribution function of treatment effect $F_{Y_{1}-Y_{0}}^{\ast}(\delta)$ (Example 2.2), which corresponds to the choice $\varphi(y_{1},y_{0})=y_{1}-y_{0}$. As a special case, the fraction of positive treatment effect (Example 2.1) is given by $\mathbb{P}(Y_{1}>Y_{0})=1-F_{Y_{1}-Y_{0}}^{\ast}(0)$.
@@
where $\tilde{\varphi}_{y}(\delta|x)=\sup\left\{y_{0}\in\mathcal{Y}_{0}(x):\varphi(y,y_{0})<\delta\right\}$. For a fixed $\delta$, the set $\left\{y_{0}\in\mathcal{Y}_{0}(x):\varphi(y,y_{0})<\delta\right\}$ may be empty for some $y\in\mathcal{Y}_{1}(x)$ and $x\in\mathcal{X}$. In such cases, we define $\tilde{\varphi}_{y}(\delta|x)$ as minus infinity.
Fan et al. (2017) show that when $\psi$ is a $\varphi$-indicator function, the identified set $\widetilde{\Theta}_{I}$ based on $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$ corresponds to the interval $\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$, where $F_{\varphi}^{L}(\delta)=\mathbb{E}[F_{\min,\varphi}(\delta|X)]$ and $F_{\varphi}^{U}(\delta)=\mathbb{E}[F_{\max,\varphi}(\delta|X)]$. Proposition 3.1 therefore implies that the identified set $\Theta_{I}$ under the experimental data is also given by $\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$. Fan et al. (2017) also examine the identification power of the covariates $X$ through these analytical expressions for the sharp bounds.
We can similarly characterize the identified set $\widetilde{\Theta}_{IC}$ based on $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$. Let $\mathcal{Y}_{1}(s,x)$ and $\mathcal{Y}_{0}(s,x)$ denote the supports of $Y_{1}$ and $Y_{0}$ given $(S,X)=(s,x)$, respectively. Then $\widetilde{\Theta}_{IC}$ corresponds to the interval $\left[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)\right]$, where $F_{L,\varphi}(\delta)=\mathbb{E}[F_{\min,\varphi}(\delta|S,X)]$ and $F_{U,\varphi}(\delta)=\mathbb{E}[F_{\max,\varphi}(\delta|S,X)]$ with
@@
Therefore, by Theorem 3.3, the identified set $\Theta_{IC}$ under the combined data is given by the interval $\left[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)\right]$ when $\theta_{o}$ is represented by a $\varphi$-indicator function. The key difference from the experimental-data case lies in the inclusion of the self-selection variable $S$ in equations (12) and (13). Since $S$ may encode information about the dependence structure between $Y_{1}$ and $Y_{0}$, its inclusion can lead to a tighter identified set.
@@
Proposition 4.3 (Identified Sets for $\varphi$-Indicator Functions).
@@
Proposition 4.3(ii) is our novel contribution, deriving the sharp bounds for $\theta_{o}$—when represented by a $\varphi$-indicator function—under the combined experimental and observational data. Together with Lemma 3.2, this result enables direct computation of the identified set from the combined data.
@@
To illustrate the identifying power of the combined data, we present simple numerical examples in Appendix A. Focusing on $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$, these examples demonstrate that the combined data can substantially narrow the identified set—and may even yield point identification—when self-selection depends on the potential outcomes.
Similar to Theorem 4.2, for a $\varphi$-indicator function we can establish sufficient and necessary conditions under which $\Theta_{IC}$ is a proper subset of $\Theta_{I}$. To simplify the technical argument, the following theorem presents this result for the case where $\mathcal{Y}_{d}(s,x)=\mathcal{Y}_{d}$ for $d=0,1$ and all $(s,x)\in\{0,1\}\times\mathcal{X}$.
Theorem 4.4 (Identification Power of Data Combination).
Suppose that $\psi$ is a $\varphi$-indicator function with $\varphi$ being continuous and non-decreasing in each argument, and that the condition in Proposition 4.3(ii) holds. Suppose further that $\mathcal{Y}_{d}(s,x)=\mathcal{Y}_{d}$ for $d=0,1$ and all $(s,x)\in\{0,1\}\times\mathcal{X}$. Then $\Theta_{IC}=\Theta_{I}$ if and only if, for almost all $x\in\mathcal{X}$, there exist $\bar{y}(x)$ and $\underline{y}(x)$ such that, for both $s\in\{0,1\}$, the function
$y\mapsto[F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1]$ (14)
attains its maximum at $\bar{y}(x)$ and its minimum at $\underline{y}(x)$.
The condition in Theorem 4.4 requires that the locations at which the function (14) attains its maximum and minimum be invariant to the self-selection variable $S$. This invariance condition is easily violated when self-selection $S$ depends on the potential outcomes $(Y_{1},Y_{0})$, as outcome-dependent selection typically alters the relative ordering of the conditional distributions across different values of $S$. In such cases, the theorem implies that combining experimental and observational data tightens the identified set.
A sufficient condition for the “if and only if” statement in Theorem 4.4 is that the self-selection variable $S$ is independent of the potential outcomes $(Y_{1},Y_{0})$ conditional on $X$. Under this selection-on-observables assumption, combining the two data sources provides no additional identifying power.
@@ 5 Additional Restrictions and Computational Approach
As in the philosophy of partial identification analysis (Manski, 2003), additional model restrictions can, when plausible, further narrow the identified set for $\theta_{o}$. Such restrictions, however, may complicate the analysis, particularly by making it difficult to derive sharp bounds. In this section, we introduce two restrictions that are plausible in many empirical contexts and present a computational approach to obtain the identified set under these restrictions with combined data.
@@ 5.1 Additional Restrictions
The first restriction is a form of positive dependence between the potential outcomes, specifically the mutually left-tail decreasing (LTD) condition (Joe, 2014). The potential outcomes $Y_{1}$ and $Y_{0}$ are said to be mutually LTD if they satisfy the following condition.
Assumption 5.1 (Positive Dependence).
The conditional distributions $\mathbb{P}(Y_{1}\leq t\mid Y_{0}\leq y,S=s,X=x)$ and $\mathbb{P}(Y_{0}\leq t\mid Y_{1}\leq y,S=s,X=x)$ are each non-increasing in $y$ almost everywhere, for almost all $(s,x)\in\{0,1\}\times\mathcal{X}$. [^11] Conditioning on $S$ and $X$ may render this assumption restrictive. Alternatively, one may impose the same condition without conditioning on $S$ and $X$, resulting in a weaker assumption with less identifying power.
This assumption implies that individuals with higher potential outcomes under one treatment state are more likely to have higher potential outcomes under the other state. Such an assumption is plausible in many empirical contexts. For example, in a small-class-size program, students who perform well academically in either small- or regular-sized classes are also likely to perform well in the alternative class size. Frandsen and Lefgren (2021) consider a slightly stronger assumption and demonstrate that it can substantially tighten the identified set for the distribution of treatment effects. [^12] Frandsen and Lefgren (2021) consider a mutually stochastically increasing assumption, which requires the conditional distributions $\mathbb{P}(Y_{1}\leq t\mid Y_{0}=y,S=s,X=x)$ and $\mathbb{P}(Y_{0}\leq t\mid Y_{1}=y,S=s,X=x)$ to be non-increasing in $y$. This assumption is stronger than Assumption 5.1 (see Joe, 1997, Theorem 2.3). The same or related assumptions are also employed by Chetty et al. (2017) in their empirical study of income mobility and by Cui and Han (2025) in policy learning with distributional welfare.
@@
This assumption implies that individuals who self-select into treatment are more likely to experience higher treatment effects than those who self-select into no-treatment. We refer to this as the generalized Roy model selection assumption, as it is implied by the selection behavior in the generalized Roy model. Because this restriction pertains to the self-selection mechanism, it cannot be leveraged using experimental data alone. Thus, an additional benefit of incorporating observational data is the ability to exploit such behavioral restrictions on self-selection.
@@ 5.2 Computational Approach
We propose a computational approach to obtain the sharp bounds for $\theta_{o}$ under the combined data, incorporating either or both Assumptions 5.1 and 5.2. When these assumptions are imposed in addition to Assumptions 2.12.5, the sharp lower and upper bounds for $\theta_{o}$ can be obtained by solving the following minimization and maximization problems:
@@
$F_{Y_{d}|Y_{d^{\prime}}\leq y,S=s,X=x}(t)\geq F_{Y_{d}|Y_{d^{\prime}}\leq y^{\prime},S=s,X=x}(t)$ (17)
[…]
@@
The constraint in (16) follows from Theorem 3.3, which shows that the identified set for $\theta_{o}$ is fully characterized by $F_{Y_{1}SX}^{\ast}$ and $F_{Y_{0}SX}^{\ast}$. The constraints in (17) and (18) are directly derived from Assumptions 5.1 and 5.2, respectively. Note that the optimization problem (15)–(18) is formulated without using the treatment variable $D$ or the data source indicator $G$, since the constraint in (16) already incorporates all identifying information conveyed by these variables (Theorem 3.3). This formulation reduces the computational burden by minimizing the number of optimization variables.
When all random variables are discrete, the optimization problem (15)–(18) reduces to a finite-dimensional linear program, for which efficient algorithms and solvers are available. [^14] Many empirical applications, however, involve continuous outcomes and covariates. A common practical approach is to discretize these continuous variables, though this may come at the potential cost of losing sharpness in identification. The details of this linear programming formulation are provided in Appendix B. Inference methods for linear program solutions, including those proposed by Fang et al. (2023) and Cho and Russell (2024), are applicable in this setting.
Let $\widetilde{\mathcal{F}}^{\dagger}$ denote the class of CDFs $F$ for all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5 and 5.15.2 with $F^{\ast}$ replaced by $F$. The identified set for $\theta_{o}$ under the combined data and these assumptions is defined as
$\Theta_{IC}^{\dagger}\equiv\big\{\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]:F\in\widetilde{\mathcal{F}}^{\ast}\big\}$,
@@
The following proposition shows that $\Theta_{IC}^{\dagger}$ corresponds to an interval whose lower and upper bounds are given by the solutions to the minimization and maximization problems (15)–(18).
@@
This result enables us to compute the identified set for $\theta_{o}$ under the additional restrictions (Assumptions 5.1 and 5.2) by solving the linear program (15)–(18). Although our analysis focuses on the case in which both assumptions are imposed jointly, sharp bounds can also be obtained from optimization problems of the same form when each assumption is imposed individually.
@@ 6 Empirical Illustration
We illustrate the proposed approach using data from the DRPT study of Gaines and Kuklinski (2011), collected in Illinois as part of the 2008 Cooperative Campaign Analysis Project. The study examines the effect of negative campaign advertisements on candidate evaluations during the 2008 U.S. presidential election.
The sample consists of 483 adult Illinois residents who were randomly assigned to one of three groups: treatment ($n=118$), no-treatment ($n=129$), and self-selection ($n=236$). Individuals in the treatment group were exposed to negative campaign advertisements (e-flyers) about John McCain and Barack Obama, whereas those in the no-treatment group were not. Participants in the self-selection group were allowed to choose whether to view the advertisements, with 90 out of 236 participants opting in.
The outcome variable $Y$ is each respondent’s feeling thermometer rating toward each candidate, measured on a scale from 0 (very unfavorable) to 100 (very favorable), with 50 indicating neutrality. The dataset includes a single categorical covariate, $X$, indicating respondents’ partisanship: Republican ($n=207$), Democrat ($n=233$), and Independent ($n=43$).
Let $Y_{d,\text{McCain}}$ and $Y_{d,\text{Obama}}$ denote the potential feeling thermometer ratings toward John McCain and Barack Obama, respectively, under treatment status $d\in\{0,1\}$, where $d=1$ indicates exposure to the negative campaign advertisements. Our parameter of interest is $\mathbb{P}(Y_{1,j}<Y_{0,j})$, which represents the proportion of individuals whose evaluation of candidate $j\in\{\text{McCain},\text{Obama}\}$ is negatively affected by the campaign material. This parameter is particularly appealing because it remains well-defined and substantively meaningful even when the outcome $Y_{d,j}$ is ordinal rather than cardinal. This feature is important in our application, as the feeling thermometer rating reflects individuals’ subjective assessments that may lack a consistent cardinal interpretation (see, e.g., Wilcox et al. (1989) for discussion). In such a context, common causal parameters, such as the ATE $\mathbb{E}[Y_{1,j}-Y_{0,j}]$, may fail to provide a meaningful interpretation.
For each candidate $j\in\{\text{McCain},\text{Obama}\}$, we estimate the identified set of $\mathbb{P}(Y_{1,j}<Y_{0,j})$ both with and without inclusion of the self-selection sample ($G=\mathrm{obs}$) and with and without imposing Assumption 5.1 (Positive Dependence). Assumption 5.1 is motivated by the idea that individuals who hold a more favorable view of a candidate in the control state are likely to maintain relatively favorable views even when exposed to negative campaign advertisements, inducing positive dependence between the potential outcomes. We do not impose Assumption 5.2, as the self-selection of negative campaign advertisements is unlikely to follow the Generalized Roy selection model. Without Assumption 5.1, the identified sets are estimated using Proposition 4.3 (with $\varphi(y_{1},y_{0})=y_{1}-y_{0}$ and $\delta=0$), together with Lemma 3.2, replacing $F_{YDGX}^{\ast}$ with its empirical distribution. When Assumption 5.1 is imposed, the identified sets are estimated via linear programming using empirical distributions (see Appendix B for details).
@@
Table 1 reports the estimated identified sets of $\mathbb{P}(Y_{1}<Y_{0})$ for each candidate, for the full population and specified subpopulations (Democrats, Republicans, and each self-selection type $s\in\{0,1\}$), across the four combinations of (i) including vs. excluding the self-selection sample ($G=\mathrm{obs}$) and (ii) imposing vs. not imposing Assumption 5.1. The estimated identified sets based solely on the random-assignment sample ($G=\mathrm{exp}$) are wide and therefore not particularly informative. By contrast, incorporating the self-selection sample markedly tightens the identified sets—by 22% and 79% for John McCain without and with Assumption 5.1, respectively, and by 24% and 67% for Barack Obama without and with Assumption 5.1.
In particular, the joint use of the self-selection sample and Assumption 5.1 (our baseline setup) yields sufficiently informative estimates. For John McCain, the baseline results suggest that at least 40% of individuals are negatively affected by the negative campaign advertisement, while at least 47% appear resistant. Qualitatively similar patterns are observed for Barack Obama (see Table 1). These findings point to a substantial yet highly heterogeneous impact of the negative advertisements.
Overall, these findings demonstrate the empirical value of combining random-assignment and self-selection data sources. They highlight a novel advantage of DRPT designs: improving the informativeness of DTE analyses through the incorporation of self-selection data.